3/8 of light initial intensity is absorbed in the beaker with length of 2cm filled with 9%- solution. Find the length of the beaker should be used to get transmittance 1/3 for the same solution.
1
Expert's answer
2020-10-25T18:32:15-0400

According to Beer-Lambert law for both cases:


"\frac{I_1}{I_0}=e^{-kl_1}=\frac{3}{8}.\\\space\\
\frac{I_2}{I_0}=e^{-kl_2}=\frac{1}{3}."

Our goal is to find l2:


"\frac{e^{l_2}}{e_{0.02}}=\frac{9}{8},\\
l_2=13.7\text{ cm}."